OpenStudy (anonymous):

Write the equation of the line given by the following graph in standard form. 3x – y = –3 3x – y = 3 –3x + y = –3 –3x – y = 3

OpenStudy (zzr0ck3r):

we know the slope is -2 and a point is (-4,0) so y = -2x+b so 0 = -2(-4) + b so b = -8 y = -2x-8

OpenStudy (campbell_st):

the line passes through the points (-4, 0) and (-2, -4) then the slope of the line is m = ( -4 -0)/(-2 - -4) m = -4/2 m = -2 the slope of the line is m = -2 so non of the listed answers fit the line

OpenStudy (anonymous):

how can I simplify this y = -2x-8

OpenStudy (campbell_st):

you can't

OpenStudy (anonymous):